A population has a mean of 300 and a standard deviation of 60. Suppose a sample of size 100 is selected and is used to estimate . Use z-table.

What is the probability that the sample mean will be within +/- 5 of the population mean (to 4 decimals)?

What is the probability that the sample mean will be within +/- 12 of the population mean (to 4 decimals)?
Answer #1

Mean, = 300

Standard deviation, = 60

Sample size, n = 100

Sample mean, = 300

Standard error, =

=

= 6

P(sample mean will be within +/- 5 of the population mean) = P(-5/ < Z < 5/)

= P(-5/6 < Z < 5/6)

= P(-0.83 < Z < 0.83)

= P(Z < 0.83) - P(Z < -0.83)

= 0.7969 - 0.2033

= 0.5936

P(sample mean will be within +/- 12 of the population mean) = P(-12/ < Z < 12/)

= P(-12/6 < Z < 12/6)

= P(-2 < Z < 2)

= P(Z < 2) - P(Z < -2)

= 0.9772 - 0.0228

= 0.9544